OpenStudy (anonymous):

Lecture #8 homework check. Most of A&E chapter 8 problems do not have answers in the back of the book, so I would like to swap answers for those problems.

OpenStudy (anonymous):

\[\text{A&E 8-16) }Cr^{3+}Cl^-_3 \text{ has lower lattice energy than }Cr^{2+}Cl^-_2\]\[\text{A&E 8-18) }Mg^{2+}\text{ }Ba^{2+}\text{ }Fe^{3+}\text{ }U \propto \frac{z_+}{r_+}\]A&E 8-22) No, the Lewis dot symbol is a bookkeeping model. It does not address polarity or geometry. A&E 8-27) S; N; B; S\[\text{A&E 8-34) By using }sp^3d^x \text{ hybrid orbitals. This must result in lower energy.}\]A&E 8-58) Exothermic. SrO forms a very stable ionic bond with oxygen. The valance structure of the atoms reflects the relative charges of the ions.\[\text{A&E 8-60) }H_f(CaO(s))=H_f(Ca(g))+H_f(O(g))+IE(Ca)+EA(O)+U\]\[U=H_f(CaO(s))-(H_f(Ca(g))+H_f(O(g))+IE_{1+2}(Ca)+EA_{1+2}(O))\]\[-3500\frac{kJ}{mol}\approx -635.09-(177.8+249.2+(589.8+1145.4)+(-142+844))\]\[H_{cryst}=-3527\frac{kJ}{mol}\text{ is the value I found on the net, so this looks right.}\]\[\text{A&E 8-74) ICl looks reasonable; } NO_4^- \text{ seems goofy}\]\[\text{A&E 8-80) }H-C\equiv N: \hspace{24pt}\text{| No reasonable resonance structures.}\]A&E 8-90) I can't draw these without the lewis packages, so I'll paste links to diagrams: http://wiki.answers.com/Q/What_is_the_lewis_structure_for_icl3 http://www.cod.edu/dept/chem/poc/problhtm/lewisstructures-clderivatives/lewisstructure/ls-38.html http://en.wikipedia.org/wiki/Thionyl_chloride http://bilbo.chm.uri.edu/CHM401/fall2006exam1-4.html